The maximum value of y=5x2 - 2x +1 is?? Explanation

Grade:11

1 Answers

Pooman Arora
2 years ago
Equation 5x2 -2x +1 will have maximum value when 
Dy/dx =0 
So dy/dx of 5x2 -2x +1 = 10x-2 
10x - 2 =0 
So x = 1 /5 will be the maximum value of equation above